OpenStudy (anonymous):

Triangle ABC is congruent to triangle DEF. In triangle ABC, side AB measures 4, side BC measures x+8, and side CA measures 8. In triangle DEF, side DE measures 4, side EF measures 3x+3, and side FD measures 8. What equation would help you to solve for the side length of BC and EF? Explain your reasoning using complete sentences.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@wolf1728 plz help

OpenStudy (welshfella):

side BC is congruent to side EF so x + 8 = 3x + 3

OpenStudy (welshfella):

solve this equation for x then you you can calculate the length of the sides BC and EF

OpenStudy (anonymous):

@welshfella can u help me??

OpenStudy (anonymous):

@welshfella I got 5/2 is that correct

OpenStudy (welshfella):

x + 8 = 3x + 3 subtract x from both sides 8 = 2x + 3 add 3 to both sides 11 = 2x divide both sides by 2 x = 5.5

OpenStudy (welshfella):

oops i made an error.. 8 = 2x + 3 subtract 3 from both sides 5 = 2x x = 2.5

OpenStudy (welshfella):

or 5/2 thats the value of x now you have to plug this into x + 8 to find the length of BC and EF.

OpenStudy (welshfella):

BC = 2.5 + 8

OpenStudy (welshfella):

ok?

OpenStudy (anonymous):

ok thank you very much (: